1. Classify the following numbers as rational or irrational:

(i) 2 - 5 (ii) (3 + 23) − 23
(iii) 2 7 (iv) 1
7 7 2

(v) 2π

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i) 2 - 5 is an irrational number, as difference of a rational and an irrational number is irrational. (ii) (3+ 23)− 23 =3+ 23− 23 =3, 8 is a rational number. (iii) 72 77 = 72 , is a rational number. (iv) 12 = 22 of an irrational number by a non-zero rational number is irrational. (v) 2π , irrational number, as π is an irrational number and multiplication of a rational and an irrational number is irrational.